OpenStudy (anonymous):

CAn i get help wth these three cal 1 problems

OpenStudy (dumbcow):

yes you can :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

anyone can help

OpenStudy (dumbcow):

for #1: \[\int\limits_{b}^{a}f(x) =-\int\limits_{a}^{b}\] \[\int\limits_{2}^{6}f(x) =\int\limits_{2}^{4}f(x)+\int\limits_{4}^{6}f(x)\] \[\int\limits\limits_{2}^{4}f(x) =\int\limits\limits_{0}^{4}f(x)-\int\limits\limits_{0}^{2}f(x\]

OpenStudy (unklerhaukus):

\[\int\limits_6^2f(x)\text dx =-\int\limits_2^6 f(x)\text dx\]\[=\int\limits_0^4f(x)\text d x-\int\limits_0^2f(x)\text d x+\int\limits_4^6f(x)\text dx\]\[=-1-3+5\]\[=1\]\[=a\]

OpenStudy (anonymous):

would the answer be -1

OpenStudy (dumbcow):

yes

OpenStudy (dumbcow):

for #3 \[\int\limits_{1}^{5}f(x)+2g(x) = \int\limits_{1}^{5}f(x)+2\int\limits_{1}^{5}g(x) = (F(5)-F(1))+2(G(5)-G(1))\]

OpenStudy (dumbcow):

remember the fundamental thm of calculus